### v1

# Controlling Ensemble Variance in Diffusion Models: An Application for Reanalyses Downscaling

Abstract
In recent years, diffusion models have emerged as powerful tools for generating ensemble members in meteorology. In this work, we demonstrate that a Denoising Diffusion Implicit Model (DDIM) can effectively control ensemble variance by varying the number of diffusion steps. Introducing a theoretical framework, we relate diffusion steps to the variance expressed by the reverse diffusion process. Focusing on reanalysis downscaling, we propose an ensemble diffusion model for the full ERA5-to-CERRA domain, generating variance-calibrated ensemble members for wind speed at full spatial and temporal resolution. Our method aligns global mean variance with a reference ensemble dataset and ensures spatial variance is distributed in accordance with observed meteorological variability. Additionally, we address the lack of ensemble information in the CARRA dataset, showcasing the utility of our approach for efficient, high-resolution ensemble generation.
keywords: Diffusion Models, Ensemble Generation, Reanalysis Downscaling, Calibrating Variance
1 Introduction
Neural approaches to downscaling meteorological reanalyses have proven effective, enhancing spatial and temporal resolutions while reducing computational costs compared to traditional numerical models.
The task is typically framed as an image-to-image prediction, transforming a low-resolution input into a high-resolution output. Widely adopted prediction-based models, such as U-Net [1] and ViT [2], operate deterministically and predict the most likely outcome for a given input. While this can produce high-quality predictions, these models lack the ability to quantify uncertainty or capture extreme events. To address this limitation, conditioned generative approaches, such as diffusion models [3], have gained traction in the scientific community. Unlike deterministic models, diffusion models sample outcomes from a learned distribution, producing diverse ensemble members even when conditioned on the same low-resolution input. These ensembles provide valuable insights into uncertainty estimation, the underlying physical system, and extreme events.
In this paper, we apply a Denoising Diffusion Implicit Model (DDIM) [4] to downscale meteorological reanalyses, focusing on wind speed. Specifically, we perform the transformation from a global reanalysis (ERA5 [5]) to a regional reanalysis (CERRA [6]) over the European domain. Building on prior work [7], we develop a diffusion model capable of operating across the full CERRA domain and we explore architectural modifications to scale diffusion models to large domains. A critical aspect of generative diffusion is the statistical properties of the generated distribution, particularly the variance. We aim to mathematically analyze how variance relates to the model’s architecture and ensure its spatial distribution aligns with physical expectations. Our analysis begins with a mathematical exploration of the reverse diffusion process, revealing that the variance of ensemble members relates to the number of diffusion steps. We experimentally validate this relationship in the ERA5-to-CERRA task, leveraging the CERRA-EDA [6] ensemble dataset for comparison. Our results demonstrate that adjusting the number of diffusion steps provides sufficient control to match the variance expressed by CERRA-EDA. This establishes our method as a practical tool for tuning ensemble diffusion to align with known uncertainty. Finally, studying the spatial distribution of the variance, we provide a methodology to obtain the optimal number of diffusion steps in specific applications, and we showcase our approach by applying it to a dataset of the Artic region lacking ensemble information, CARRA [8].
1.1 Background
Reanalysis datasets are a cornerstone of modern climate and weather research, providing gridded, high-resolution reconstructions of past atmospheric conditions. Reanalysis are widely used by researchers, institutions, and organizations to support studies ranging from climate variability to extreme event prediction [9, 10, 11]. High-resolution reanalyses are often created using downscaling techniques, which refine coarse global climate data into finer spatial and temporal scales. Traditional dynamical downscaling is computationally expensive, requiring significantly more resources than coarser global climate models and leading to data gaps and incomplete assessments of model uncertainty and regional climate variability [12, 13].
In recent years, deep learning has emerged as an effective alternative for downscaling, offering powerful and cost-effective methods for producing high-resolution data [14, 15]. A variety of deep learning models for downscaling have been proposed [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], showcasing their ability to capture complex spatial patterns and generate high-quality outputs across a range of meteorological variables.
Among these methods, diffusion models stand out for their ability to capture the uncertainty of the downscaling process, enabling not only the generation of high-resolution data but also produce probabilistic ensembles by sampling from a latent space. This capability makes them highly effective for modeling uncertainty and variability, which are crucial in meteorological applications [29, 30, 31, 32, 33, 34, 35].
Downscaled ensemble datasets are essential for quantifying uncertainties [36, 37, 38], detecting extreme events [39], and performing ensemble-based statistical analyses critical for assessing the likelihood of compound extreme events [40, 41]. In recent years an increasing amount of publications focused on the utilization of generative deep learning architectures in the downscaling task, utilizing the inherently probabilistic nature of the models to generate ensemble members.
In [20], the authors extend generative adversarial networks (GANs) to downscale precipitation forecasts, addressing the challenges of adding fine-scale structure and correcting forecast errors. Using high-resolution radar data as ground truth, the study demonstrates that GANs and VAE-GANs can produce spatially coherent precipitation maps while maintaining statistical properties comparable to state-of-the-art post-processing methods. In [32], the authors propose using deep generative diffusion models to emulate and enhance probabilistic weather forecasting. Learning from historical data, diffusion model generate realistic ensemble forecasts at a fraction of the computational cost of traditional physics-based simulations. The generated ensembles replicate the statistical properties and predictive skill of operational forecasts, while also correcting biases, improving reliability, and better capturing probabilities of extreme weather events. This scalable approach has potential applications in both weather forecasting and large-scale climate risk assessment. In [18], the authors propose a statistical downscaling method for satellite sea surface wind (SSW) data using generative adversarial networks (GANs) with a dual learning scheme. The dual learning approach combines high-resolution SSW reconstruction with degradation kernel estimation in a closed loop. In [38], the authors propose a cascading deep learning framework based on generative diffusion models for downscaling tropical cyclone (TC) wind fields from low-resolution ERA5 data (0.25°) to high-resolution observations (0.05°) at 6-hour intervals. The framework consists of a debiasing neural network to improve wind speed accuracy and a conditional denoising diffusion probabilistic model (DDPM) for super-resolution and ensemble generation. Applied to 85 TC events, the model achieves realistic spatial structures, enabling accurate high-resolution TC wind modeling and risk assessment. In [37], the authors present a hybrid downscaling framework combining dynamical downscaling with generative diffusion models to enhance resolution and improve uncertainty estimates for regional climate projections. The method uses a regional climate model (RCM) to downscale Earth System Model (ESM) outputs to intermediate resolution, followed by a diffusion model for further refinement. Applied to the CMIP6 ensemble, this approach provides more accurate uncertainty bounds, lower errors than bias correction and spatial disaggregation (BCSD), and better captures spectra and multivariate correlations. The framework offers a cost-effective and accurate alternative for downscaling large climate projection ensembles.
1.2 Diffusion models basics
Diffusion models, and in particular Denoising Diffusion Implicit Models (DDIM) [4], are state-of-the-art generative models able to generate highly realistic data from noise. In particular, given a sequence $\{\alpha_{t}\}_{t=1}^{T}\subseteq(0,1)$ such that $\alpha_{0}\approx 0$ and $\alpha_{T}\approx 1$, called noise schedule, a DDIM defines an iterative procedure to turn random noise $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ into a new sample $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$, where $p_{gt}(\boldsymbol{x}_{T})$ represents the probability distribution from which the training data is sampled [^1] Note that here we use a slightly different notation compared to e.g. [4], as we consider the time running reversely. We made this choice as this notation simplifies the mathematical derivation of our formula in section 3.1.. This procedure is obtained through a simple yet effective idea: let $\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T-1},\boldsymbol{x}_{T}\}$ be a sequence such that $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$, and $\boldsymbol{x}_{t}\sim p_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{T})$ for any $t=T-1,\dots,0$. If $p_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{T})=\mathcal{N}(\sqrt{\alpha_{t}}\boldsymbol{x}_{T},(1-\alpha_{t})\boldsymbol{I})$, then:
$\boldsymbol{x}_{t}=\sqrt{\alpha_{t}}\boldsymbol{x}_{T}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t},\quad\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}).$ (1)
Note that the previous equation implies that $\boldsymbol{x}_{t}$ is obtained by interpolating real data $\boldsymbol{x}_{T}$ with random noise $\boldsymbol{\epsilon}_{t}$, with interpolation coefficients $\sqrt{\alpha_{t}}$ and $\sqrt{1-\alpha_{t}}$. Therefore, they are usually referred to as signal rate and noise rate, respectively. Please refer to section 3.1 for a detailed discussion on how these terms are selected.
Clearly, Equation (1) represents the process that slowly corrupts a true image $\boldsymbol{x}_{T}$ into pure noise $\boldsymbol{x}_{0}$. To revert this process and generate data from noise, a DDIM considers a neural network $\boldsymbol{\epsilon}_{\Theta}^{t}(\boldsymbol{x}_{t})$ trained to extract the noise component $\boldsymbol{\epsilon}_{t}$ from $\boldsymbol{x}_{t}$, i.e.
$\boldsymbol{\epsilon}_{\Theta}^{t}(\boldsymbol{x}_{t})\approx\boldsymbol{\epsilon}_{t},\quad\forall t=T-1,\dots,0.$ (2)
Given that, one can simply invert the process in Equation 1 by sampling $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and then iteratively updating:
$\boldsymbol{x}_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}\boldsymbol{x}_{t-1}+\left(\sqrt{1-\alpha_{t}}-\sqrt{\frac{\alpha_{t}(1-\alpha_{t-1})}{\alpha_{t-1}}}\right)\boldsymbol{\epsilon}_{\Theta}^{t-1}(\boldsymbol{x}_{t-1}).$ (3)
After $T$ steps, this process leads to a new sample $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$. For a detailed introduction to DDIM, please refer to [4, 42, 43].
A property of DDIM that is crucial for our work is the possibility of skipping a few steps in the generative process. This is achieved by setting a step-size $\Delta t\in\mathbb{N}$ and modifying Equation 3 as:
$\boldsymbol{x}_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\boldsymbol{x}_{t-\Delta t}+\left(\sqrt{1-\alpha_{t}}-\sqrt{\frac{\alpha_{t}(1-\alpha_{t-\Delta t})}{\alpha_{t-\Delta t}}}\right)\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}).$ (4)
Note that the choice $\Delta t=1$ recovers the original generative process, while $\Delta t>1$ allows for a faster generation, possibly at the expense of quality. This modification can be relevant when the neural network model $\boldsymbol{\epsilon}_{\Theta}^{t}(\boldsymbol{x}_{t})$ is particularly large, or when the dimensionality of the data $\boldsymbol{x}_{T}$ is big. Indeed, when this happens, computing $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})$ could be time-prohibitive, and reducing the amount of time this operation has to be done to reach the end of the process is crucial for its applicability. Since the number of iterations of Equation 4 equals to $N=\frac{T}{\Delta t}$, increasing $\Delta t$ leads to a faster generation process.
In this work we argue that the importance of the selection of the step-size $\Delta t$ is not limited to the time-efficiency of the generative process, but it can also be used to control the variance of the sampled data: a key aspect when these models are employed for generating ensembles.
1.3 Notations
Throughout this article, we will make use of some notations, which we report here for completeness. In particular, we use lower-case non-bold latin characters to indicate scalar values, lower-case bold latin characters to indicate vectors, and upper-case bold latin characters to indicate matrices. Moreover, we will denote as $\boldsymbol{0}$, $\boldsymbol{1}$ the vector of all zeros and all ones, respectively, and as $\boldsymbol{I}$ the identity matrix. If $\boldsymbol{x}$ and $\boldsymbol{y}$ are two vectors with the same dimension, then $\boldsymbol{x}\geq\boldsymbol{y}$ means that $x_{i}\geq y_{i}$ for every $i$. The data is therefore assumed to be a vector of dimension $n$, which corresponds to the flattened version of the $(S\times M\times h\times w)$-dimensional array if it is indicated with a lowercase bold letter, while it is treated as a tensor if indicated with an uppercase bold letter. For example, if $\boldsymbol{X}\in\mathbb{R}^{S\times M\times h\times w}$ represents the usual tensor-shaped dataset with $S$ datapoints, $M$ channels, and spatial dimension $h\times w$, then $\boldsymbol{x}\in\mathbb{R}^{n}$ represents its flattened version, with $n=S\cdot M\cdot h\cdot w$.
1.4 Structure of the article
The remainder of this paper is organized as follows. In Section 2 we introduce the datasets employed for the experiments, namely CERRA, CERRA-EDA and CARRA, representing a high-resolution world map with positional information of the wind. In Section 3 we provide the theoretical background for the proposed idea, and show a possible interpretation of the observed phenomena. Next we introduce the methodology applied in the successive experiments, described in Section 5. Finally, in Section 6 we conclude the paper by summarizing our results.
2 Data
In this section we describe the primary datasets used in this study. Our task involves using the global reanalyses ERA5 [44] as the low-resolution input to condition our model in predicting a high-resolution downscaled counterpart, trained on the regional reanalyses CERRA [6]. The resulting ensemble members are validated against CERRA-EDA, an ensemble version of CERRA at half the temporal and spatial resolution. We will also apply our methods to CARRA, a dataset built on the same core model as CERRA and that covers the Arctic region, but lack ensemble members. This section begins with a brief introduction to ERA5, followed by a focused discussion on CERRA, CARRA, and CERRA-EDA.
2.1 ERA5
ERA5 [5], the fifth generation ECMWF reanalysis, is a globally recognized dataset extensively utilized in climate and atmospheric research. Spanning from 1940 to the present, it provides hourly estimates of atmospheric, land, and oceanic variables at a 0.25° horizontal resolution (approximately 30 km) and 137 vertical levels. Based on the ECMWF Integrated Forecasting System, ERA5 incorporates advanced numerical weather prediction models and satellite data, ensuring high reliability and precision [44]. Its widespread use spans diverse applications, including precipitation trends, temperature analysis, wind studies, and extreme event research, making it an indispensable resource in climate science.
2.2 CERRA
The Copernicus Regional Reanalysis for Europe (CERRA [6]) is a sophisticated high-resolution Regional ReAnalysis (RRA) dataset specifically designed for the European region. It is a product of the European Copernicus Climate Change Service (C3S), executed through a contract with the Swedish Meteorological and Hydrological Institute (SMHI), in collaboration with subcontractors Meteo-France and the Norwegian Meteorological Institute. CERRA operates at a high 5.5 km horizontal resolution, offering enhanced spatial detail for meteorological variables across Europe.
(a) Domain of CERRA and CERRA-EDA
(b) Domain of CARRA-EAST
Figure 1: Comparison of CERRA and CARRA-EAST domains.
To achieve this fine resolution, CERRA relies on the global ERA5 reanalysis for initial and boundary conditions. In addition to ERA5, CERRA incorporates an extensive array of high-resolution observational data, covering conventional and satellite sources, along with physiographic datasets that represent surface characteristics. One innovative aspect of CERRA is its integration with the CERRA-EDA (Ensemble of Data Assimilation), which is a 10-member ensemble system that provides essential background error statistics. These statistics, recalculated every second day, enable CERRA to dynamically adjust initial conditions, significantly enhancing the accuracy of data assimilation and model outputs.
CERRA provides a comprehensive range of meteorological variables, including temperature, humidity, wind speed and direction, precipitation, and cloud cover, all presented on a consistent $1069\times 1069$ grid. Unlike ERA5’s hourly structure, CERRA’s temporal resolution combines reanalysis and forecast data, featuring eight reanalysis time points daily at three-hour intervals starting from 00 UTC. Each reanalysis cycle is followed by a six-hour forecast period, with additional 30-hour forecasts generated at 00 UTC and 12 UTC. For this study, we focus exclusively on the reanalysis data. CERRA employs a Lambert Conformal Conic (LCC) projection to minimize distortion at high latitudes, ensuring greater accuracy in polar regions.
2.3 CARRA
The C3S Arctic Regional Reanalysis (CARRA) dataset [8] addresses the unique environmental and climatic conditions of the Arctic region, which are insufficiently captured by global reanalyses such as ERA5. While ERA5 provides a 31 km global resolution, CARRA offers 2.5 km horizontal resolution, capturing finer details essential for understanding the Arctic’s rapidly changing climate. Observational records and climate projections show that warming in the Arctic occurs at over twice the global average rate, leading to increased environmental and economic activities and driving a need for detailed data to support climate adaptation and management in the region.
CARRA leverages the same HARMONIE-AROME weather prediction model as CERRA, enhanced for reanalysis applications at the ECMWF high-performance computing facility. It features a three-hourly analysis update with a comprehensive three-dimensional variational data assimilation scheme, incorporating diverse data sources. This includes an extensive array of local observations from Nordic countries and Greenland, advanced satellite-derived glacier albedo data, improved sea ice and snow initialization, and high-resolution physiography and orography data, specifically adjusted for the Arctic’s unique topography.
The dataset spans two key domains:
  • CARRA-West, covering Greenland, the Labrador Sea, Davis Strait, Baffin Bay, Denmark Strait, Iceland, Jan Mayen, the Greenland Sea, and Svalbard.
  • CARRA-East, which includes Svalbard, Jan Mayen, Franz Josef Land, Novaya Zemlya, Barents Sea, and the northern regions of the Norwegian Sea and Scandinavia.
These domains encompass the Arctic’s four largest land ice bodies, including the Greenland Ice Sheet and the Austfonna Ice Cap, which are focal points in climate research. CARRA’s high resolution allows for accurate representation of critical Arctic-specific weather phenomena, such as polar lows in the North Atlantic and Greenland’s downslope katabatic winds, which have significant implications for local communities and maritime activities.
The CARRA reanalysis is driven by ERA5 global reanalysis data for boundary conditions, integrating enhanced observations and ArcticDEM-based elevation data. CARRA also provides critical improvements in data quality through the assimilation of satellite observations (e.g., radiances, scatterometer data) and refined upper-air data algorithms, making it a powerful tool for Arctic climate monitoring. Covering the period from September 1990 to the present, CARRA supports comprehensive climate analysis with monthly updates that maintain a 2-3 month latency, ensuring the dataset remains relevant and accurate for ongoing climate research and operational needs.
For our works we will focus on the CARRA-EAST domain.
2.4 CERRA-EDA
CERRA-EDA [45] is an Ensemble of Data Assimilation (EDA) system comprising 10 members. It is based on the same code as the deterministic CERRA system and uses the same observational data, with the exception of the Microwave Sounding Unit and has a lower 11-km horizontal resolution.
Each ensemble member in CERRA-EDA is generated by perturbing observations prior to data assimilation. One control member is initialized without any perturbations, while the other nine members incorporate Gaussian-distributed random perturbations within the estimated observational errors. No perturbations are applied to Sea Surface Temperature (SST), sea ice, or physical parameters. The lateral boundary conditions for CERRA-EDA are provided by the ERA5 ensemble, which includes 10 members and operates with a 63-km horizontal grid.
The analyses generated by CERRA-EDA are used to produce six-hour forecasts, from which background error covariances are estimated. CERRA relies on CERRA-EDA’s output, which supplies essential background error statistics and is integral to the data assimilation process.
CERRA-EDA is run in advance of the deterministic system, though this typically does not hinder production due to CERRA-EDA’s faster processing enabled by its lower resolution and longer time steps, with the 10 ensemble members operating in parallel.
3 Methodology
In this section we quickly review our signal and noise scheduling methodology before delving into the mathematical reasoning behind the control of variance via the number of diffusion steps. In particular, we derive an explicit formula to predict the variance of the generated data, which shows how it depends on the number of diffusion steps.
3.1 Signal and noise rates
Figure 2: Reverse diffusion signal and noise ratios with cosine scheduling, including the reduced signal rate
In Section 1 we discussed the importance of the diffusion schedule $\{\alpha_{t}\}_{t=1,\dots,T}$ in the generation process of DDIM. Recently, various schedules have been proposed to find a perfect balance between the signal and noise in the diffusion process, including linear schedules [3] and sinusoidal schedules [46, 4]. In this study, we adopt sinusoidal functions due to their ability to achieve smooth and controlled transitions between signal and noise components, which is crucial for effective denoising and high-quality image generation. In particular, we consider the following choice for the signal rate as in [42]:
$\sqrt{\alpha_{t}}=\sin\left(\pi\frac{T-t}{T}\right).$ (5)
Therefore, the noise rate becomes:
$\sqrt{1-\alpha_{t}}=\cos\left(\pi\frac{T-t}{T}\right),$ (6)
for which it holds that $\left(\sqrt{\alpha_{t}}\right)^{2}+\left(\sqrt{1-\alpha_{t}}\right)^{2}=\sin\left(\pi\frac{T-t}{T}\right)^{2}+\cos\left(\pi\frac{T-t}{T}\right)^{2}=1$, ensuring that the noisy images, formed as a weighted sum of signal and noise, consistently maintain unit variance [46, 47]. The extremes of the schedule (i.e. when $t=0$ and $t=T$) can be problematic, as near-pure signal or noise levels often cause instability. Following common practice [47, 48], we clamp the extremes to maintain a balanced Signal-to-Noise Ratio (SNR), preventing degenerate cases and ensuring stable training and reliable reverse diffusion.
As for the time discretization, we consider a uniform subdivision of the $[0,T]$ range into $T$ intervals, so that $t\in\{0,1,\dots,T\}$. Note that, when a step-size $\Delta t>1$ is considered, then clearly the time-domain becomes:
$t\in\{0,\Delta t,2\Delta t,\dots,N\Delta t\},$ (7)
where $N=\frac{T}{\Delta t}$ represents the number of timesteps of our generative process.
3.2 Controlling the variance
The generative process for DDIM involves a sequence of denoising operations alternated with the re-injection of the same noise. Our findings reveal that the key aspect of controlling the variance is selecting the number of diffusion steps. In this section, we analyze the evolution of the variance through the iterative diffusion process and we provide an intuition on how the number of diffusion steps $N$ influences the variance of the generated data. We remark that the diffusion process in DDIM is defined by sampling $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and then iterating through:
$\boldsymbol{x}_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}),$ (8)
where we defined $c_{t-\Delta t}$ as:
$c_{t-\Delta t}:=\sqrt{1-\alpha_{t}}-\sqrt{\frac{\alpha_{t}(1-\alpha_{t-\Delta t})}{\alpha_{t-\Delta t}}},$ (9)
to simplify the notation. We also recall that the diffusion schedule $\{\alpha_{t}\}_{t=0,\dots,T}$ is selected such that $\alpha_{0}\approx 1$ and $\alpha_{T}\approx 0$. Note that $\alpha_{T}\neq 0$ as we clamped its value to avoid instabilities during training, as we already discussed in section 3.1.
Let $\boldsymbol{x}_{0}$ be an $n$-dimensional random variable with Gaussian distribution such that $\mathbb{E}[\boldsymbol{x}_{0}]=\boldsymbol{0}$ and $Var(\boldsymbol{x}_{0})=\boldsymbol{I}$, and let $\{\boldsymbol{x}_{t}\}_{t=0,\dots,T}$ be the stochastic process defined in (8). Since the full variance matrix $Var(\boldsymbol{x}_{t})$ has dimension $n\times n$, which is prohibitive from a computational point of view, in this work we limit our analysis to the element-wise variance, defined as:
$\boldsymbol{v}(\boldsymbol{x}_{t})_{i}:=Var(\boldsymbol{x}_{t})_{i,i},$ (10)
which is a vector of dimension $n$, representing the variance between each pixel of $\boldsymbol{x}_{t}$ with itself, ignoring the relationship with the adjacent pixels. Note that this choice is coherent with our task, as we want to study the pixel-wise uncertainty of DDIM downscaling, and taking into consideration the correlation between adjacent pixels is beyond the scope of this analysis.
It is not difficult to show (see Appendix A) that the element-wise variance has similar properties to the classical variance, and in particular:
  1. (P1)
    $\boldsymbol{v}(a\boldsymbol{x}_{t})=a^{2}\boldsymbol{v}(\boldsymbol{x}_{t})$,
  2. (P2)
    $\boldsymbol{v}(\boldsymbol{x}_{t}+\boldsymbol{y}_{t})=\boldsymbol{v}(\boldsymbol{x}_{t})+\boldsymbol{v}(\boldsymbol{y}_{t})+2\boldsymbol{v}(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$,
where $\boldsymbol{v}(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is the element-wise covariance between $\boldsymbol{x}_{t}$ and $\boldsymbol{y}_{t}$.
Proposition 3.1.
Let $\boldsymbol{m}_{t}:=\mathbb{E}[\boldsymbol{x}_{t}]$ and $\boldsymbol{v}_{t}:=\boldsymbol{v}(\boldsymbol{x}_{t})$. Then, for any $t=\Delta t,\dots,N\Delta t$, it holds:
$\boldsymbol{v}_{t}\approx\boldsymbol{F}_{t}\boldsymbol{v}_{t-\Delta t}+\boldsymbol{g}_{t},$ (11)
where:
$\boldsymbol{F}_{t}:=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{I}+\boldsymbol{J}_{t},$
$\boldsymbol{J}_{t}:=2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}diag\left(\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\right),$
$\boldsymbol{g}_{t}:=c_{t-\Delta t}^{2}\boldsymbol{1}.$
Proof.
See Appendix A. ∎
Note that (11) represents a first-order linear recurrence equation of dimension $n$ with non-constant coefficients, for which there exists a closed formula that can be used to compute $\boldsymbol{v}_{T}$, the variance of the generated data, as summarized in the following Theorem.
Theorem 3.2.
Let $\boldsymbol{F}_{t}$, $\boldsymbol{J}_{t}$, and $\boldsymbol{g}_{t}$ defined as in Proposition 3.1, and let $\Delta t$ be a fixed step-size for DDIM, so that any timestep $t$ can be written as $i\Delta t$ for $i=0,\dots,N$. Note that $\boldsymbol{v}_{0}=\boldsymbol{1}$, the vector of all ones, as $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. Then, defining $\boldsymbol{F}_{(N+1)\Delta t}$ to be the identity matrix to simplify the notation, it holds:
$\boldsymbol{v}_{T}\approx\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)\boldsymbol{1}+\sum_{i=1}^{N}\left[\left(\prod_{k=i+1}^{N+1}\boldsymbol{F}_{k\Delta t}\right)\boldsymbol{g}_{i\Delta t}\right].$ (12)
Proof.
See Appendix A. ∎
Notably, Theorem 3.2 not only allows for the explicit computation of the element-wise variance of $\boldsymbol{x}_{T}$, but it also underlines the dependence of $\boldsymbol{v}_{T}$ on the number of diffusion timesteps. Indeed, if, for example, both $\boldsymbol{F}_{i\Delta t}\geq\boldsymbol{0}$ and $\boldsymbol{g}_{i\Delta t}\geq\boldsymbol{0}$, then Equation 12 clearly implies that $\boldsymbol{v}_{T}$ monotonically grows as a function of $N$. Specifically, increasing the step size $\Delta t$ (and consequently reducing the number of timesteps $N$) results in fewer terms being included in the summation that computes $\boldsymbol{v}_{T}$, thereby decreasing its value.
4 Experimental Setting
In this section, we describe the pre-processing, training, and evaluation strategies required to adapt the diffusion model for high-resolution downscaling across the full CERRA domain. The transition to larger spatial scales necessitated adjustments to both data handling and model architecture to ensure consistent performance. We focus on data selection and pre-processing in Section 4.1, followed by a detailed explanation of the modifications required for training on the full domain in Section 4.2.
4.1 Data selection and pre-processing
The diffusion model performs downscaling in an image-to-image manner, transforming low-resolution ERA5 data into high-resolution CERRA or CARRA outputs while maintaining alignment over the same spatial domain for input and output. The primary experimental framework of our study focuses on downscaling wind speed from ERA5 to CERRA over the full CERRA domain and comparing the ensemble results with the CERRA-EDA dataset. Our model will be trained solely on the CERRA data and no input of CERRA-EDA is used during training, ensuring that the model will learn the ensemble variability only from the CERRA data itself.
The CERRA domain spans a grid of 1069×1069 points across Europe, represented in a Lambert Conformal Conic (LCC) projection. The complete CERRA dataset covers the period from September 1984 to June 2021; for our experiments, we utilized data from 1985 to 2010 for training and from 2016 for ensemble testing. Additionally, we evaluated overall performance for the entire decade from 2011 to 2020, though computational constraints limited ensemble variance testing to the year 2016. In our diffusion model, ERA5 data serves as the conditioning input, re-gridded to align with CERRA’s LCC grid. Wind speed, $s$, is derived from the two wind components, $s_{u}$ and $s_{v}$, as:
$s:=\left\|\begin{bmatrix}s_{u}\\ s_{v}\end{bmatrix}\right\|_{2}=\sqrt{s_{u}^{2}+s_{v}^{2}}$
Our diffusion model is based on a residual convolutional denoiser, which benefits from input dimensions divisible by two at multiple levels. To ensure consistency between the downsample and upsample paths, we augment the original grid size (1069×1069) with mirroring padding, resulting in a final input size of 1072×1072. Our conditioning is implemented via concatenation on the input of the denoising network, to equalize the dimension we opt to upscale the ERA5 data to $1069\times 1069$ via bilinear interpolation, and then we apply the same mirroring padding to ensure spatial consistency.
To compare our diffusion-based ensembles with the existing CERRA-EDA ensemble members, we note that CERRA-EDA operates at half the spatial resolution of CERRA (565×565) and a temporal resolution of six hours, also halved relative to CERRA. For consistency, we train our model exclusively on full-resolution CERRA data with a three-hour temporal resolution. Wind speed is an instantaneous value, therefore six-hourly outputs can be generated using conditioning information at the same temporal resolution. The generated outputs retain the full CERRA resolution; for comparison with CERRA-EDA, we apply bilinear interpolation to downscale the results.
Diffusion models require standardized inputs to balance signal and noise rates effectively. We apply a standardizer that normalizes inputs using the mean and variance computed across the training set, ensuring consistent scaling across batches. It is important to note that this variance is relative to the training data and is distinct from the ensemble variance between members produced by the diffusion model. This normalization is applied to both high-resolution images (signal rates) and low-resolution conditioning data, aligning their scales and simplifying the learning task. Experimentally, this approach demonstrated improved performance and stability.
4.2 Training on the full domain
The general structure of our diffusion model follows the framework established in [7]. The model utilizes a residual U-Net [1, 49] as a denoiser. Training is conducted by uniformly sampling signal-to-noise ratios from the sinusoidal noise scheduling functions, optimizing the mean absolute error (MAE) between the sampled noise and the predicted noise.
Scaling the diffusion model to the full CERRA domain, with image sizes of 1069×1069, required modifications to a model originally designed for downscaling images up to 256×256. These adaptations involved adjustments to both the noise schedule and the network architecture to ensure effective performance at higher resolutions.
Recent findings [47, 48] have demonstrated that cosine noise schedules can reduce the effective noise added by the model as image size increases. This occurs because redundancy in data, such as correlations among nearby pixels, typically increases with resolution, and the independently added noise becomes easier to denoise in larger images. Consequently, noise schedules optimized for smaller resolutions may not perform as effectively at higher resolutions. To address this issue, we introduce a constant scaling factor $\lambda>1$ to reduce the signal rate in the noise schedule. For an input $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$ and a scaling factor $\lambda$, the noise schedule eq. 1 is modified as follows:
$\boldsymbol{x}_{t}=\sqrt{\alpha_{t}}\frac{\boldsymbol{x}_{T}}{\lambda}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t}$ (13)
Experimentally, setting the value $\lambda=3$ improved training efficiency and overall model performance, which is a crucial aspect for large-scale ensemble diffusion applications. A comparison of signal and noise ratios with the scaled signal rate is shown in Figure 2. Notably, applying the same scaling to the low-resolution conditioning signal further stabilized the training process.
While the core U-Net architecture remained largely unchanged, scaling to larger domains required increasing the number of residual blocks in the bottleneck. This adjustment improved performance when generating larger images [48]. The model’s performance across the increased domain remained comparable to that observed at smaller resolutions, with the bottleneck modification enhancing its ability to process higher-resolution inputs effectively.
The diffusion model is trained on the ERA5 to CERRA task for wind speed on the full domain, utilizing data from 1985 to 2010 for training. The training is run on three A100 GPUs, utilizing the AdamW optimizer.
5 Experiments
Our primary experimental objective is to evaluate the behavior of ensemble variance while changing the timestep $\Delta t$. While a comprehensive analysis of the ensemble diffusion model’s performance across the entire CERRA domain is beyond the scope of this paper, we ensure that the model operates within a realm of satisfactory performance. To achieve this, we begin—similar to [7]—by comparing our model’s performance against bilinearly interpolated ERA5 data and a residual U-Net model trained for the same task. The results, obtained by computing the Mean Squared Error (MSE) and SSIM between the high-resolution output produced by the model and the original CERRA, are shown in Table 1.
Bilinear U-Net DDIM (Our)
[…]
SSIM ()($\uparrow$)( ↑ ) $0.751$ $0.892$ $0.923$
Table 1: Performance of ensemble diffusion against bilinear interpolation and U-Net for the years 2011-2020.
Testing for the years 2011–2020, the ensemble diffusion model achieves an MSE of $2.54\cdot 10^{-4}$, compared to $2.50\cdot 10^{-4}$ for the U-Net and $3.73\cdot 10^{-3}$ for bilinear interpolation. For SSIM, bilinear interpolation scores 0.751, the U-Net scores 0.892, and ensemble diffusion achieves the highest score of 0.923. Ensemble diffusion was executed with $N=2$ diffusion steps and an ensemble of 10 members.
These results confirm that the diffusion model delivers satisfactory performance, with MSE values approximately an order of magnitude better than bilinear interpolation and comparable to those of the U-Net. Additionally, as expected, the diffusion model outperforms both alternatives in terms of SSIM. The proposed metrics were calculated on normalized data.
5.1 Quantifying variance
Figure 3: Plotting the global mean variance $\mu_{V}$ of ensemble diffusion for different number of steps against CERRA-EDA
The first aspect of variance analysis we evaluate in our experiments is understanding how the model’s expressed variance changes with the number of diffusion steps, as detailed in Section 3.2. For this experiment we selected the year 2016 and computed the ensemble variance for 10 members for all the year at 6h intervals, matching the existing format of the CERRA-EDA ensemble. Furthermore, we used bilinear interpolation to reduce the spatial dimension of our output image from $1069\times 1069$ to $565\times 565$, also matching the existing CERRA-EDA format. Having obtained a comparable dataset, we proceeded to calculate the pixel-wise variance for the ensemble member of each ensemble of the year, finally calculating the mean variance, a value representing the overall variance expressed by the diffusion model.
Given a dataset of ensembles $\boldsymbol{D}\in\mathbb{R}^{S\times M\times h\times w}$, where $S$ represents the number of samples for 2016, $M=10$ is the number of ensemble members for each datapoint, and $(h,w)=(565,565)$ are the spatial dimensions, we compute the pixel-wise variance across the ensemble members $\boldsymbol{V}\in\mathbb{R}^{S\times M\times 565\times 565}$, such that for each sample $i=1,\dots,S$, and any spatial coordinate $(x,y)\in\{1,\dots,h\}\times\{1,\dots,w\}$,
$\boldsymbol{V}_{i}(x,y)=\frac{1}{M}\sum_{j=1}^{M}\left(\boldsymbol{D}_{i,j}(x,y)-\frac{1}{M}\sum_{k=1}^{M}\boldsymbol{D}_{i,k}(x,y)\right)^{2}.$ (14)
We then compute the mean variance across all samples and spatial locations as:
$\mu_{\boldsymbol{V}}=\frac{1}{S\cdot h\cdot w}\sum_{i=1}^{S}\sum_{x=1}^{h}\sum_{y=1}^{w}\boldsymbol{V}_{i}(x,y).$ (15)
Here, $\mu_{\boldsymbol{V}}$ is a scalar value representing the global mean variance across all samples and pixels. By computing $\mu_{\boldsymbol{V}}$ for different diffusion steps, we observe that the global variance increases with the number of steps, as illustrated in Figure 3. Furthermore, we demonstrate that the variance control mechanism successfully aligns the global variance with that of CERRA-EDA, with the closest match achieved at $N=12$ diffusion steps.
5.2 Analyzing variance spatially
Figure 4: Comparing the spatial variance of CERRA-EDA with ensemble diffusion at different diffusion steps. The comparison relative to the testing year 2016 and is performed in a 3 monthly manner to highlight seasonal variations.
The global mean-variance inform us that the diffusion model is able to match the overall amount of variance present in CERRA-EDA, but this gives us no information to where the variance is distributed in space. Our interest here is to examine if there is a match between the two spatial distributions of the variance.
For the following experiments, we decided to investigate the variance in a three-monthly setting, to highlight the behavior of the model in different seasons. To achieve this goal we divided the the samples into four sets, including respectively January February and March (JFM), April May and June (AMJ), July August and September (JAS) and October November and December (OND).
To analyze spatial variance, we compute the pixel-wise mean of variance maps across all samples within each three-monthly subset. Let $k\in\{JFM,AMJ,JAS,OND\}$ denote the season, $\mathcal{S}_{k}\subseteq\{1,\dots,S\}$ represent the subset of indices corresponding to the season $k$, and $S_{k}$ the number of samples corresponding to season $k$. The spatial mean variance $\mu^{spatial}_{k}$ for season $k$ is given by:
$\boldsymbol{\mu}^{spatial}_{k}(x,y)=\frac{1}{S_{k}}\sum_{i\in\mathcal{S}_{k}}\boldsymbol{V}_{i}(x,y),\quad(x,y)\in\{1,\dots,h\}\times\{1,\dots,w\}.$ (16)
Here, $\boldsymbol{V}_{i}$ and $\boldsymbol{\mu}^{spatial}_{k}$ are understood to represent spatial fields, where the value of each pixel is computed independently. We compute the spatial mean-variance for the reference ensemble CERRA-EDA and for Ensemble Diffusion across a range of diffusion steps $N\in\{2,4,8,10,12,14\}$ for the year 2016. This allows us to visualize how the spatial distribution of variance evolves as the number of diffusion steps increases. The obtained results are reported in Figure 4. We observe substantial overlap between the variance distributions of CERRA-EDA and Ensemble Diffusion. This indicates that the diffusion model successfully captures the variability of the data by training solely on the CERRA dataset, without requiring additional information about the underlying uncertainty. Furthermore, we note that the degree of spatial overlap varies by season, with the highest overlap observed in the JFM and OND sets, while AMJ and JAS exhibit less overlap.
To numerically quantify the overlap between variance distributions, we introduce the mean-variance discrepancy (MVD) score, a metric defined as the absolute error between the spatial mean-variance maps for each season. Let $\boldsymbol{\mu}_{k,\text{ED}}^{\text{spatial}}$ denote the spatial mean-variance map of Ensemble Diffusion and $\boldsymbol{\mu}_{k,\text{CE}}^{\text{spatial}}$ denote the corresponding map for CERRA-EDA. The MVD score for season $k$ is then defined as:
$\text{MVD}_{k}=\frac{1}{h\cdot w}\sum_{x=1}^{h}\sum_{y=1}^{w}\left|\boldsymbol{\mu}_{k,\text{ED}}^{\text{spatial}}(x,y)-\boldsymbol{\mu}_{k,\text{CE}}^{\text{spatial}}(x,y)\right|.$ (17)
The numerical results for the MVD scores computed across all three-monthly periods are reported in Table 2.
From the MVD results, we observe that the optimal yearly number of diffusion steps is $N=8$, a lower value if compared with 12, which was identified based on the global mean variance. This finding reveals that, while 12 steps yield an overall variance closest to the reference, the spatial distribution of the variance behaves differently. This underscores the importance of a tailored tuning procedure to achieve better alignment with expected uncertainties. Moreover, the MVD scores are not uniform throughout the year: the winter months achieve the lowest MVD scores with 4 diffusion steps, while the summer months perform best with 12 diffusion steps, possibly due to the model being affected by seasonal dynamics.
Alongside the full-domain analysis, we examine the behavior of the generated ensemble at a randomly selected point within the domain to provide a clearer and more readable visualization of the mean and variance ranges. For this analysis, we plot the mean value and variance range at the chosen point for both the 8-step ensemble diffusion and CERRA-EDA. To improve clarity over the high-frequency time series, we apply a moving average smoothing with a 30-sample window. The resulting plot, shown in Figure 5, reveals consistent overlap between the mean and variance ranges.
MVD Scores for Seasonal Spatial Variance Maps
2 steps 4 steps 8 steps 10 steps 12 steps 14 steps
[…]
Table 2: Mean variance discrepancy (MVD) scores computed between the spatial mean variance maps of Ensemble Diffusion and CERRA-EDA for each three-monthly period (JFM, AMJ, JAS, OND) and each timestamp for 2016. Lower scores indicate higher similarity between the variance distributions of the two ensembles.
5.3 Impact of $\Delta t$ on performance
An interesting aspect of the variance analyses is to check how the overall performance of the Ensemble Diffusion (ED) model changes as the number of steps increase. To achieve this goal we have have utilized $M=10$ ensemble members at $565\times 565$ spatial resolution and 6-hourly time resolution for the year 2016. We have computed the ensemble diffusion mean and compared it with the available CERRA-EDA control run, which is, as mentioned in Section 2.4, the member of the ensemble with no perturbation on the model input. We evaluate the comparison via MSE, calculated on normalized values. The results, reported in table 3, reveal that the performance is largely stable while changing the step size $\Delta t$, with a slight increase in the performance for larger step sizes. This should be expected as more steps allows the reverse diffusion process to make smaller, more gradual corrections at each step, leading to a more precise reconstruction of the target high-resolution image.
MSE ensemble mean performance comparison
2 steps 4 steps 8 steps 10 steps 12 steps 14 steps
[…]
Table 3: Table comparing the MSE performance of the diffusion model at different numbers of reverse diffusion steps. The MSE is calculated between the mean of the 10 generated ensembles and the CERRA-EDA control run. The results are reported 3-monthly for the year 2016.
Figure 5: Mean and variance range at a randomly selected point in the domain, compared between ensemble diffusion (8 reverse diffusion steps) and CERRA-EDA. Data is smoothed using a 30 samples moving average window for readability and covers the entire year of 2016.
5.4 Demonstrating capabilities on CARRA
Both CERRA and CARRA, introduced in Section 2 are built upon the HARMONIE-AROME weather prediction system [50], adapted for reanalysis purposes. This common foundation ensures that both systems share a core methodology for atmospheric modeling and data assimilation, making them methodologically similar despite their focus on different regions.
Having validated our diffusion model’s capability to generate ensemble members for the CERRA dataset, we naturally extend this methodology to CARRA-EAST, the Arctic region dataset focusing on the eastern domain. Unlike CERRA, CARRA lacks an ensemble dataset, which limits its usability in quantifying variability, validating against real-world uncertainties, detecting rare or extreme events, performing ensemble-based statistical analyses, and estimating model robustness. Our ensemble diffusion driven ensemble dataset can be a highly efficient way to address this issues.
For CARRA-EAST, we adopt a training setup similar to the one proposed for CERRA. The downscaling process is framed as an image-to-image transformation, converting low-resolution ERA5 data into high-resolution CARRA-EAST, mantaining the same spatial domain between the two, with the resulting grid having a size of $989\times 789$. We regrid ERA5 wind speed data to match the projection of CARRA-EAST. The training period spans 1991–2010, while the year 2016 is reserved for ensemble testing. As with CERRA, padding is applied to adjust the grid size for compatibility with the diffusion architecture. Training was conducted following the setup detailed in Section 4.2.
The ensemble members are generated for the year 2016 using 8 diffusion steps, the optimal value determined through spatial validation with MVD on CERRA-EDA. Figure 6 illustrates the spatial variance for 2016, plotted within the same variance range as Figure 4. Our results reveal that the spatial variance is consistent with expected patterns.
Figure 6: Revealing the spatial variance of ensemble diffusion trained on CARRA-EAST with 8 diffusion steps. The comparison is relative to the testing year 2016 and is performed in a 3 monthly manner to highlight seasonal variations.
6 Conclusions
[…]
Since our diffusion model is trained solely on the deterministic ERA5-to-CERRA task, it is unaffected by factors such as measurement quality or other previously assessed uncertainties. As such, our approach can serve as a validation tool for existing ensembles, with discrepancies providing insights into the neural model’s understanding of the underlying physical system or serving as validation for uncertainty estimates obtained through other methods. In our experimental setting, the spatial variance reported in Figure 4 shows a clear overlap between distributions. However, during the summer months in the North African region, discrepancies between CERRA-EDA and our Ensemble Diffusion appear. This may indicate that our model struggles to capture uncertainty in this region, or that CERRA-EDA itself may be affected by data scarcity [6].
In conclusion, our methodology provides a practical tool for tuning ensemble diffusion models to produce the correct variance distribution. We anticipate that future implementations will consider the number of diffusion steps as a crucial tuning parameter. Furthermore, our theoretical framework opens promising directions for studying the statistical properties of ensemble members in relation to the structure of diffusion models, encouraging further research.
7 Acknowledgements
ERA5, CERRA, CERRA-EDA and CARRA were downloaded from the Copernicus Climate Change Service (C3S) (2023). This research was partially funded and supported by the following Projects:
  • European Cordis Project “Optimal High Resolution Earth System Models for Exploring Future Climate Changes” (OptimESM), Grant agreement ID: 101081193
  • ISCRA Project “AI for weather analysis and forecast” (AIWAF)
Statements and Declarations
The authors declare no competing interests.
Code availability
The code and downscaled ensemble members relative to the presented work is archived at the following GitHub https://github.com/fmerizzi/ERA5-to-CERRA-via-Diffusion-Models/. We also provide the full list of the environmental libraries and respective versions.
Appendix A Proofs of main results
In this section, we report the derivation of the main results presented in the paper.
A.1 Properties of element-wise variance
At first, we need to show that element-wise variance $\boldsymbol{v}(\boldsymbol{x}_{t})$, defined in (10) inheritates similar properties as the full variance, and in particular (P1) and (P2). This is not hard to prove, as:
$\boldsymbol{v}(a\boldsymbol{x}_{t})_{i}=Var(a\boldsymbol{x}_{t})_{i,i}=a^{2}Var(\boldsymbol{x}_{t})_{i,i}=a^{2}\boldsymbol{v}(\boldsymbol{x}_{t})_{i},$
and:
$\boldsymbol{v}(\boldsymbol{x}_{t}+\boldsymbol{y}_{t})_{i}$ $=Var(\boldsymbol{x}_{t}+\boldsymbol{y}_{t})_{i,i}=Var(\boldsymbol{x}_{t})_{i,i}+Var(\boldsymbol{y}_{t})_{i,i}+2Cov(\boldsymbol{x}_{t},\boldsymbol{y}_{t})_{i,i}$
$=\boldsymbol{v}(\boldsymbol{x}_{t})_{i}+\boldsymbol{v}(\boldsymbol{y}_{t})_{i}+2\boldsymbol{v}(\boldsymbol{x}_{t},\boldsymbol{y}_{t})_{i},$
which proves the required properties.
A.2 Proof of Proposition 3.1
The essence of our theoretical results lies on the recursive formula introduced in Proposition 3.1, which follows by properties (P1) and (P2) applied to the diffusion process (4). In particular note that, by (4),
$\boldsymbol{v}(\boldsymbol{x}_{t})=\boldsymbol{v}\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\right).$
Then, applying (P1) and (P2),
$\boldsymbol{v}(\boldsymbol{x}_{t})=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+$ $c_{t-\Delta t}^{2}\boldsymbol{v}(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))$
$+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})).$
Since by (2), $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\approx\boldsymbol{\epsilon}_{t-\Delta t}$, then $\boldsymbol{v}(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))\approx\boldsymbol{v}(\boldsymbol{\epsilon}_{t-\Delta t})=\boldsymbol{1}$, which results in:
$\boldsymbol{v}(\boldsymbol{x}_{t})=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+c_{t-\Delta t}^{2}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})).$
To simplify this expression, recall the definition of the covariance:
$\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))=\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})-\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})])^{T}\right],$
where we denoted $\boldsymbol{m}_{t-\Delta t}:=\mathbb{E}\left[\boldsymbol{x}_{t-\Delta t}\right]$, and consider the first order Taylor decomposition of $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})$ around $\boldsymbol{m}_{t-\Delta t}$:
$\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\approx\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})+\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right).$
Consequently:
$\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})]$ $\approx\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})+\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)]$
$=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})+\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\mathbb{E}\left[\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)\right]$
$=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t}),$
where in the first equality we used the linearity of the expected value and the observation that $\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})]=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})$ as $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})$ is constant with respect to $\boldsymbol{x}_{t-\Delta t}$, while the second equality follows by observing that:
$\mathbb{E}\left[\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)\right]=\mathbb{E}[\boldsymbol{x}_{t-\Delta t}]-\mathbb{E}[\boldsymbol{m}_{t-\Delta t}]=\boldsymbol{m}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}=\boldsymbol{0}.$
Plugging the equations above into the definition of covariance, we get:
$\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))$ $\approx\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})-\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t}))^{T}\right]$
$\approx\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right))^{T}\right]$
$=\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right))^{T}\right]$
$=\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t}).$
Substituting this result back into the equation for the variance of $\boldsymbol{x}_{t}$, we get:
$\boldsymbol{v}(\boldsymbol{x}_{t})$ $=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+c_{t-\Delta t}^{2}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})$
$=\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\right)\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+c_{t-\Delta t}^{2}\boldsymbol{1},$
which corresponds to Proposition 3.1 by setting:
$\boldsymbol{F}_{t}:=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t}),$
$\boldsymbol{g}_{t}:=c_{t-\Delta t}^{2}\boldsymbol{1}.$
A.3 Proof of Theorem 3.2
The proof of the next Theorem is mainly a recursive application of Proposition 3.1 to compute $\boldsymbol{v}_{T}$ from $\boldsymbol{v}_{0}$. Indeed, we recall that:
$v_{t}\approx\boldsymbol{F}_{t}\boldsymbol{v}_{t-\Delta t}+\boldsymbol{g}_{t}.$
Since, as already observed, $\boldsymbol{v}_{0}=\boldsymbol{1}$, then:
$\boldsymbol{v}_{\Delta t}\approx\boldsymbol{F}_{\Delta t}\boldsymbol{1}+\boldsymbol{g}_{\Delta t},$
$\boldsymbol{v}_{2\Delta t}\approx\boldsymbol{F}_{2\Delta t}\boldsymbol{F}_{\Delta t}\boldsymbol{1}+\boldsymbol{F}_{2\Delta t}\boldsymbol{g}_{\Delta t}+\boldsymbol{g}_{2\Delta t},$
$\boldsymbol{v}_{3\Delta t}\approx\boldsymbol{F}_{3\Delta t}\boldsymbol{F}_{2\Delta t}\boldsymbol{F}_{\Delta t}\boldsymbol{1}+\boldsymbol{F}_{3\Delta t}\boldsymbol{F}_{2\Delta t}\boldsymbol{g}_{\Delta t}+\boldsymbol{F}_{3\Delta t}\boldsymbol{g}_{2\Delta t}+\boldsymbol{g}_{3\Delta t},$
$\dots$
$\boldsymbol{v}_{N\Delta t}\approx\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)\boldsymbol{1}+\sum_{i=1}^{N}\left(\prod_{k=i+1}^{N+1}\boldsymbol{F_{k\Delta t}}\right)\boldsymbol{g}_{i\Delta t},$
where we defined $\boldsymbol{F}_{(N+1)\Delta t}:=\boldsymbol{I}$ for simplicity, concluding the proof.